# Conditional mutual information: A generalization of causal inference in quantum systems

### v1

@@ Abstract
The concept of causality is fundamental to numerous scientific explanations; nonetheless, its extension to the quantum regime has yet to be explored rigorously. This paper introduces the development of a quantum causal index, a novel extension of the classical causal inference framework, tailored to grasp the causal relationships inherent in quantum systems. Our study focuses on the asymmetric quantum conditional mutual information (QCMI), incorporating the von Neumann entropy, as a directional metric of causal influence in quantum many-body systems. We analyze spin chains using the QCMI, implementing a projective measurement on one site as the intervention and monitoring its effect on a distant site conditioned on intermediate spins. Additionally, we study the effective causal propagation velocity, which is the speed at which QCMI becomes significant at distant sites. These findings indicate the presence of finite-speed propagation of causal influence, along with the emergence of coherent oscillations.
I Introduction
The ‘cause-effect’ relationship between the variables of dynamical systems remains one of the fundamental questions in the literature for a long time [1, 2, 3, 4]. Establishing a causal inference, which distinguishes correlation from actual influence, is essential for developing more accurate models and effective interventions. Therefore, understanding causal relationships is crucial for studying dynamical systems, as it clarifies how dynamic states evolve and how various components of a system interact. A fundamental assumption in causal reasoning is based on the ‘Reichenbach Principle of Common Cause’, which philosopher Hans Reichenbach originally articulated in his seminal work, The Direction of Time, published in $1956$ [5]. Following this principle, if $X$ and $Y$ are two statistically correlated variables and neither $X$ causes $Y$ nor $Y$ causes $X$, then there exists a third variable $Z$, referred to as a common cause, that explains this correlation.
@@ I Introduction
In $2001$, Paluš et al. [14] proposed a specific form of conditional mutual information (CMI), $I(A;B|C)$, which is asymmetric in variables $A$ and $B$ and can be used to measure the direction of information flow from $A$ to $B$ when the third variable $C$ is given. The notion of this particular form of CMI is solely based on classical dynamics. Generally, in information theory [15], the standard form of CMI $I(A:B|C)$ is often used to measure the dependency between two variables $A$ and $B$ given the knowledge of the other variable $C$, and $I(A:B|C)$ is symmetric in $A$ and $B$. It is essential to clarify that if we intend to utilize CMI as a causal index, it must exhibit asymmetry in its relationship with variables $A$ and $B$. This requirement arises from the fundamental nature of causal indices, which are inherently asymmetric, as they seek to identify and illustrate directional relationships between variables [1]. In other words, a causal index is designed to uncover directed dependencies between two variables, often by conditioning on third variables or leveraging temporal information. The causal index $I(A;B|C)$, proposed by Paluš et al. [14], maintains this asymmetry in $A$ and $B$. Later, Paluš and Vejmelka [16] reported an updated and higher-dimensional version of CMI, which is a frequently used form of CMI in dynamics and causality studies. A brief outline of this particular form of CMI is discussed in Appendix A. Existing literature [17, 18, 19] supports that CMI $I(A;B|C)$ has been widely explored in the study of causal structure. Motivated by this asymmetric form of CMI, in this paper, we intend to extend it to study causal relations in quantum systems. More explicitly, we are interested in reporting a quantum version of CMI, $I(A;B|C)$, which exhibits asymmetry in variables $A$ and $B$ and is suitable for studying causal structure in quantum systems.
The paper is organized as follows: The detailed formalism of the quantum conditional mutual information (QCMI) is discussed in Sec. II. Subsequently, we discuss the advantage of using asymmetric QCMI over the symmetric QCMI using a GHZ-like state in Sec. III. After that, we use QCMI as a causal index to study the spin chains and elaborately discuss it in Sec. IV. Finally, we conclude the main findings in Sec. V.
II Quantum conditional mutual information (QCMI)
@@
where $\rho_{B}^{x}=\mathrm{Tr}_{C}(\rho_{BC}^{x})$, and it is called the partial trace over the subsystem $C$. Likewise, $\rho_{C}^{x}=\mathrm{Tr}_{B}(\rho_{BC}^{x})$. This quantity $I(A;B|C)_{\rho,\mathcal{M}}$ captures how much information about $B$ and $C$ is retained after a measurement on $A$, thus encoding a directionality from $A$ to $B$ mediated through $C$. In contrast to Eq. 1, Eq. II demonstrates asymmetry between the variables $A$ and $B$. As a result, we have adopted a different notation, $I(A;B|C)$, in Eq. II — rather than using $I(A:B|C)$.
III Advantage of using asymmetric QCMI for Causal Analysis
@@
IV Application of Asymmetric QCMI for Causal Analysis in Spin Chains
Quantum spin chains consist of a sequence of spins (qubits or higher-dimensional spins) arranged linearly, governed by some Hamiltonian. Such systems are key models for exploring many-body quantum physics, entanglement dynamics, and information propagation [21, 22]. Consider a spin chain with sites labeled $1,2,\ldots,N$ and each site is described by Hilbert space $\mathcal{H}_{i}$. We further define three subsystems: $A$ represents spin at site $i$, $B$ represents spin at site $j$, and $C$ is some intermediate spins. Intuitively, by applying local operations on one site (site $i$) and measuring the resulting information shared with the other site (site $j$) conditioned on intermediate subsystems (sites between $i$ and $j$), we can obtain a quantitative, directional measure of quantum causal influence that respects the dynamical constraints of many-body quantum systems.
@@ IV Application of Asymmetric QCMI for Causal Analysis in Spin Chains
Figure 1: The asymmetric QCMI $I(A;B|C)$ has been calculated for the $3$-qubit spin chain, and it is plotted as a function of time ($t$) . The corresponding parameters are: $N=3$ and $J=h=1$ in Eq. 23.
@@
Figure 2: The asymmetric QCMI $I(A;B|C)$ has been plotted as a function of time ($t$) for the four-site XX spin chain. The corresponding parameters are: $N=4$, $J=1$, and $h=0.5$ in Eq. 24.
@@
The LR bound makes the above intuition robust. It guarantees no signal can propagate faster than that exponential light-cone. The LR bound gives an exponential suppression of the effect of a local perturbation outside a causal cone of slope $v_{\mathrm{LR}}$, where $v_{\mathrm{LR}}$ is the LR velocity. Therefore, QCMI remains exponentially small outside that cone and only becomes significant when $t$ is large enough that the exponential factor ceases to be tiny. Hence, $v_{\mathrm{LR}}$ is a theoretical upper bound on propagation speed. It depends only on the Hamiltonian (coupling strengths and geometric locality) and is state-independent. Appendix B describes analytically and concisely how to calculate the LR velocity for the XX chain (Eq. 24). For the parameter choice used in the XX chain, $J=1$, we can evaluate the LR velocity as follows:
@@
Figure 3: The arrival time $t_{\mathrm{arr}}(d)$ has been plotted as a function of distance $d$. The effective causal propagation velocity $v_{\mathrm{eff}}$ can be calculated from the slope of this plot using Eq. 29. The associated parameters in Eq. 24 are: $N=8$, $J=1$, and $h=0.3$.
@@
$v(k)$ $=$ $\frac{\partial\varepsilon(k)}{\partial k},$ (32)
@@ V Conclusion
As a final point, this paper has significantly contributed to this emerging field by proposing a generalized causal index for quantum systems. It discusses the theoretical foundations of this index and emphasizes the applicability of investigating causal measures in quantum many-body systems.
@@ Acknowledgment
The author sincerely thanks Dr. M. Paluš for several fruitful discussions on classical conditional mutual information in the context of causal analysis. The author also like to extend his sincere appreciation to Dr. S. Das and Dr. S. Sur for their valuable feedbacks. Their insightful comments have significantly enhanced the quality of the manuscript.
@@ Appendix A Classical Conditional Mutual Information (CCMI)
where $H(\cdot)$ denotes the Shannon entropy [15] of the random variables, and the subscript ‘$c$’ implies that the analysis is restricted to the classical domain.
@@
where $A=x_{j}$, $B=x^{\prime}_{j+\tau}$, and $C=\{x^{\prime}_{j},x^{\prime}_{j-\eta_{0}},x^{\prime}_{j-2\eta_{0}}\}$. Additionally, $\tau$ (a positive scalar) represents time step; hence, $x^{\prime}_{j+\tau}$ is the future state. The other parameter $\eta_{0}$ is the embedding delay associated with $\{x^{\prime}_{j}\}_{j=1}^{N}$. In contrast to Eq. 35, Eq. 36 demonstrates asymmetry between the variables $X$ and $Y$. As a result, we have adopted a different notation, $I_{c}(X;Y|Z)$, in Eq. 36 — rather than using $I_{c}(X:Y|Z)$. Intuitively, this particular form of CCMI (Eq. 36) is asymmetric in $X$ and $Y$, and it quantifies the amount of information transferred from the variable $x_{j}$ to the future value $x^{\prime}_{j+\tau}$, conditioned on the present and past values of $x^{\prime}$. Note that Eq. 36 assumes well-defined time-indexed random variables drawn from observable sequences or probability distributions and is typically used in classical settings, such as time-delay embedding or information flow between coupled systems.
Appendix B Lieb–Robinson velocity for the XX chain
@@
and the single–site field term is $h_{i}=h\,\sigma_{z}^{(i)}$. We need operator–norm bounds for these terms. Using $\|\sigma_{\alpha}\|=1$ and submultiplicativity,
@@
The on–site field satisfies $\|h_{i}\|=|h|\,\|\sigma_{z}\|=|h|$, but single-site terms do not generate propagation between distinct sites and therefore do not increase the effective LR speed beyond the contribution from the two-site couplings.

### v2

@@ Abstract
The concept of causality is fundamental to numerous scientific explanations; however, its extension to the quantum regime has yet to be rigorously explored. This letter introduces the development of a quantum causal index, a novel extension of the classical causal inference framework, tailored to learn the causal relationships inherent in quantum systems. Our study focuses on the asymmetric quantum conditional mutual information (QCMI), incorporating the von Neumann entropy, as a directional metric of causal influence in quantum many-body systems. We analyze spin chains using the QCMI, implementing a projective measurement on one site as the intervention and monitoring its effect on a distant site conditioned on intermediate spins. Additionally, we study the effective causal propagation velocity, which is the speed at which QCMI becomes significant at distant sites. These findings indicate the presence of finite-speed propagation of causal influence, along with the emergence of coherent oscillations.
keywords:
Complex systems , Causal relation, Conditional mutual information , Quantum systems
[…]
[label1]organization=Department of Complex Systems,addressline=Institute of Computer Science of the Czech Academy of Sciences, city=Prague, postcode=18200, country=Czech Republic
1 Introduction
The ‘cause-effect’ relationship between the variables of dynamical systems remains one of the fundamental questions in the literature for a long time [1, 2, 3, 4]. Establishing a causal inference, which distinguishes correlation from actual influence, is essential for developing more accurate models and effective interventions. Therefore, understanding causal relationships is crucial for studying dynamical systems, as it clarifies how dynamic states evolve and how various components of a system interact. A fundamental assumption in causal reasoning is based on the ‘Reichenbach Principle of Common Cause’, which philosopher Hans Reichenbach originally articulated in his seminal work, The Direction of Time, published in $1956$ [5]. Following this principle, if $X$ and $Y$ are two statistically correlated variables and neither $X$ causes $Y$ nor $Y$ causes $X$, then there exists a third variable $Z$, referred to as the common cause, that explains this correlation.
@@ 1 Introduction
In $2001$, Paluš et al. [14] proposed a specific form of conditional mutual information (CMI), $I_{c}(A;B|C)$, which is asymmetric in variables $A$ and $B$ and can be used to measure the direction of information flow from $A$ to $B$ when the third variable $C$ is given. The subscript ‘$c$’ in $I_{c}(A;B|C)$ implies that the discussion associated with $I_{c}(A;B|C)$ is restricted to the classical domain. Generally, in information theory [15], the standard form of CMI $I_{c}(A:B|C)$ is used to measure the dependency between two variables $A$ and $B$ given the knowledge of the other variable $C$, and $I_{c}(A:B|C)$ is symmetric in $A$ and $B$. It is essential to clarify that if we intend to utilize CMI as a causal index, it must exhibit asymmetry in its relationship with variables $A$ and $B$. This requirement arises from the fundamental nature of causal indices, which are inherently asymmetric, as they seek to identify and illustrate directional relationships between variables [1]. In other words, a causal index is designed to uncover directed dependencies between two variables, often by conditioning on third variables or leveraging temporal information. The causal index $I_{c}(A;B|C)$, proposed by Paluš et al. [14], maintains this asymmetry in $A$ and $B$. Later, Paluš and Vejmelka [16] reported an updated and higher-dimensional version of CMI, which is a frequently used form of CMI in dynamics and causality studies. A brief outline of this particular form of CMI is discussed in A. Existing literature [17, 18, 19] supports that CMI $I_{c}(A;B|C)$ has been widely explored in the study of causal structure. Motivated by this asymmetric form of CMI, in this letter, we intend to extend it to study causal relations in quantum systems. More explicitly, we are interested in reporting a quantum version of CMI, $I(A;B|C)$, which exhibits asymmetry in variables $A$ and $B$ and is suitable for studying causal structure in quantum systems.
The letter is organized as follows: The detailed formalism of the quantum conditional mutual information (QCMI) is discussed in Sec. 2. Subsequently, we discuss the advantage of using asymmetric QCMI over the symmetric QCMI using a GHZ-like state in Sec. 3. After that, we use QCMI as a causal index to study the spin chains and elaborately discuss it in Sec. 4. Finally, we conclude the main findings in Sec. 5.
2 Quantum conditional mutual information (QCMI)
@@
where $\rho_{B}^{x}=\mathrm{Tr}_{C}(\rho_{BC}^{x})$, and it is called the partial trace over the subsystem $C$. Likewise, $\rho_{C}^{x}=\mathrm{Tr}_{B}(\rho_{BC}^{x})$. This quantity $I(A;B|C)$ captures how much information about $B$ and $C$ is retained after a measurement on $A$, thus encoding a directionality from $A$ to $B$ mediated through $C$. Although the final expression for QCMI presented in Eq. 2 appears formally symmetric in variables $B$ and $C$, a distinct operational difference, however, arises from the specific causal question being addressed. When evaluating whether interventions on $A$ have an influence on $B$, $B$ is designated as the target variable, while $C$ serves as the conditioning subsystem. Conversely, to investigate the influence from $A$ to $C$, one would compute $I(A;C|B)$. In contrast to Eq. 1, Eq. 2 demonstrates asymmetry between the variables $A$ and $B$. As a result, we have adopted a different notation, $I(A;B|C)$, in Eq. 2 — rather than using $I(A:B|C)$.
In determining the relationship between QCMI and CCMI (An overview of the CCMI can be found in A), we recall the following explicit form of CCMI used in classical causal inference:
which quantifies the predictive information flow from a variable $x_{j}$ to a future variable $x^{\prime}_{j+\tau}$ beyond what is contained in the past history of $B$. A positive value of $I_{c}(A;B|C)$ indicates a directed causal influence from $A$ to $B$ conditioned on $C$.
Intuitively, to get the analogy, both quantities $I_{c}(A;B|C)$ and $I(A;B|C)$ measure directed dependence from $A$ to $B$ with conditioning on past variables $\{x^{\prime}_{j},x^{\prime}_{j-\eta_{0}},x^{\prime}_{j-2\eta_{0}}\}$ corresponding to conditioning on subsystem $C$. Temporal asymmetry in the classical case is replaced by measurement back-action in the quantum case. If $\rho_{ABC}$ is diagonal in a classical basis and $\mathcal{M}_{A}$ measures in that basis,
$I(A;B|C)\longrightarrow I_{c}(x_{j};x^{\prime}_{j+\tau}\mid\{x^{\prime}_{j},x^{\prime}_{j-\eta_{0}},x^{\prime}_{j-2\eta_{0}}\}),$ (9)
so the QCMI reduces exactly to the CCMI. Detailed steps to reach CCMI from QCMI have been incorporated in B. Hence, the proposed QCMI provides a natural quantum generalization of classical information flow measures used in causal inference for dynamical systems.
3 Advantage of using asymmetric QCMI for Causal Analysis
@@
4 Application of Asymmetric QCMI for Causal Analysis in Spin Chains
Quantum spin chains consist of a sequence of spins (qubits or higher-dimensional spins) arranged linearly, governed by some Hamiltonian. Such systems are key models for exploring many-body quantum physics, entanglement dynamics, and information propagation [21]. Consider a spin chain with sites labeled $1,2,\ldots,N$ and each site is described by Hilbert space $\mathcal{H}_{i}$. We further define three subsystems: $A$ represents spin at site $i$, $B$ represents spin at site $j$, and $C$ is some intermediate spins. Intuitively, by applying local operations on one site (site $i$) and measuring the resulting information shared with the other site (site $j$) conditioned on intermediate subsystems (sites between $i$ and $j$), we can obtain a quantitative, directional measure of quantum causal influence that respects the dynamical constraints of many-body quantum systems.
@@ 4 Application of Asymmetric QCMI for Causal Analysis in Spin Chains
Figure 1: The asymmetric QCMI $I(A;B|C)$ has been calculated for the $3$-qubit spin chain, and it is plotted as a function of time ($t$). The corresponding parameters are: $N=3$ and $J=h=1.0$ in Eq. 25.
@@
Figure 2: The asymmetric QCMI $I(A;B|C)$ has been plotted as a function of time ($t$) for the four-site XX spin chain. The corresponding parameters are: $N=4$, $J=1.0$, and $h=0.5$ in Eq. 26.
@@
The LR bound makes the above intuition robust. It guarantees no signal can propagate faster than that exponential light-cone. The LR bound gives an exponential suppression of the effect of a local perturbation outside a causal cone of slope $v_{\mathrm{LR}}$, where $v_{\mathrm{LR}}$ is the LR velocity. Therefore, QCMI remains exponentially small outside that cone and only becomes significant when $t$ is large enough that the exponential factor ceases to be tiny. Hence, $v_{\mathrm{LR}}$ is a theoretical upper bound on propagation speed. It depends only on the Hamiltonian (coupling strengths and geometric locality) and is state-independent. C describes analytically and concisely how to calculate the LR velocity for the XX chain (Eq. 26). For the parameter choice used in the XX chain, $J=1.0$, we can evaluate the LR velocity as follows:
@@
Figure 3: The arrival time $t_{\mathrm{arr}}(d)$ has been plotted as a function of distance $d$. The effective causal propagation velocity $v_{\mathrm{eff}}$ can be calculated from the slope of this plot using Eq. 31. The associated parameters in Eq. 26 are: $N=8$, $J=1.0$, and $h=0.3$.
@@
$v(k)$ $=$ $\frac{\partial\varepsilon(k)}{\partial k}$ (34)
@@ 5 Conclusion
As a final point, this letter has significantly contributed to this emerging field by proposing a generalized causal index for quantum systems. It discusses the theoretical foundations of this index and emphasizes the applicability of investigating causal measures in quantum many-body systems.
@@ Acknowledgment
The author sincerely thanks Dr. M. Paluš for several fruitful discussions on classical conditional mutual information in the context of causal analysis. The author would also like to extend his sincere appreciation to Dr. N. Ganguly, Dr. S. Das, and Dr. S. Sur for their valuable comments, which have significantly enhanced the quality of this letter.
@@ Appendix A Classical Conditional Mutual Information (CCMI)
where $H(\cdot)$ denotes the Shannon entropy [15, 24, 25] of the random variables, and the subscript ‘$c$’ implies that the analysis is restricted to the classical domain.
@@
where $A=x_{j}$, $B=x^{\prime}_{j+\tau}$, and $C=\{x^{\prime}_{j},x^{\prime}_{j-\eta_{0}},x^{\prime}_{j-2\eta_{0}}\}$. Additionally, $\tau$ (a positive scalar) represents time step; hence, $x^{\prime}_{j+\tau}$ is the future state. The other parameter $\eta_{0}$ is the embedding delay associated with $\{x^{\prime}_{j}\}_{j=1}^{N}$. In contrast to Eq. 37, Eq. 38 demonstrates asymmetry between the variables $A$ and $B$. As a result, we have adopted a different notation, $I_{c}(A;B|C)$, in Eq. 38 — rather than using $I_{c}(A:B|C)$. Intuitively, this particular form of CCMI (Eq. 38) is asymmetric in $A$ and $B$, and it quantifies the amount of information transferred from the variable $x_{j}$ to the future value $x^{\prime}_{j+\tau}$, conditioned on the present and past values of $x^{\prime}$. Note that Eq. 38 assumes well-defined time-indexed random variables drawn from observable sequences or probability distributions and is typically used in classical settings, such as time-delay embedding or information flow between coupled systems.
Appendix B Classical reduction of the QCMI
[…]
$|,\rho_{ABC}=\sum_{a,b,c}p(a,b,c)\,\ket{abc}\!\bra{abc}$, (39)
[…]
$|,\rho_{BC}^{(a)}=\sum_{b,c}p(b,c|a)\,\ket{bc}\!\bra{bc}$, (40)
and their von Neumann entropies reduce to the corresponding Shannon entropies. Consequently, the asymmetric QCMI
$I(A;B|C)=\sum_{a}p(a)\!\left[S(\rho_{B}^{(a)})+S(\rho_{C}^{(a)})-S(\rho_{BC}^{(a)})\right]$ (41)
reduces exactly to the classical asymmetric CMI computed from $p(a,b,c)$,
$I(A;B|C)\longrightarrow I_{c}(x_{j};x^{\prime}_{j+\tau}\mid\{x^{\prime}_{j},x^{\prime}_{j-\eta_{0}},x^{\prime}_{j-2\eta_{0}}\}),$ (42)
where the variables $A,B,C$ correspond respectively to the source variable $x_{j}$, the future target $x^{\prime}_{j+\tau}$, and the conditioning set consisting of the past history $\{x^{\prime}_{j},x^{\prime}_{j-\eta_{0}},x^{\prime}_{j-2\eta_{0}}\}$. This limit shows that the QCMI $I(A;B|C)$ is a natural generalization of the classical asymmetric CMI used to quantify directed information flow and causality in dynamical systems.
Appendix C Lieb–Robinson velocity for the XX chain
@@
and the single-site field term is $h_{i}=h\,\sigma_{z}^{(i)}$. We need operator-norm bounds for these terms. Using $\|\sigma_{\alpha}\|=1$ and submultiplicativity,
@@
The on-site field satisfies $\|h_{i}\|=|h|\,\|\sigma_{z}\|=|h|$, but single-site terms do not generate propagation between distinct sites and therefore do not increase the effective LR speed beyond the contribution from the two-site couplings.